You are on page 1of 5

11

AUGUST 2010

TS TIDINGS

FEED BACK NO.1 PROJECT: RAICHUR THERMAL POWER STATION, UNIT 8, 250 MW PROBLEM: BOOSTER PUMP THRUST BEARING FAILURE DURING FIRST COMMISSIONING OF BFP8A. BFP 8A was started on 17.07.10 for the first time for trial run after erection clearance and checking of interlocks & protection. It tripped on BP NDE end Thrust Bearing Temperature Very High within 20 sec of pump starting. Temperatures indicated by 2 RTDs on active side BP NDE Thrust Bearing was 108.6 & 114.3 C and on non active side the temperature indicated was 52.5&57.6 C. Observation and analysis RTDs on which Pump was tripping was replaced with new one to eliminate the doubt. The orifice provided in oil supply line of BP NDE thrust Bearing was removed to increase the oil flow. Earlier orifice size was 2.5 mm. The Pump was started on 20.07.10, again tripped on BP Thrust Bearing Temperature Very High. Temperature indicated by 2 RTDs on active side was 103.1 & 107.0 C and on non active side the temperatures were 48.0 & 52.3 C. The temperature trend during Pump start up is shown in figure 1. BP NDE Drain Oil Temperature rose from 29 C to 32 C which was Normal and BP NDE Journal Bearing Temperature rose from 28 C to 32.4 C which was normal. Flow to thrust bearing was adequate, Lube oil Header Pressure was 2.6 Ksc. The float at Booster Pump was checked and was found to be 0.3 mm. Booster Pump Thrust Bearing was opened for inspection and it was found that the bearing and the thrust collar are having scorings. Photos of Thrust Collar and Bearing are enclosed for reference. The problem was referred to BHEL, Hyderabad and as per their instructions following checks were done 1. No choking / Blocking was observed in Oil supply line from header to Booster Pump NDE Thrust Bearing. 2. Orifice in BP Bearing Top housing was checked and it was through. Oil from thrust bearing flows to journal side and to drain through this hole.

TECHNICAL SERVICES / PSSR

12

AUGUST 2010

TS TIDINGS

3. Common BP NDE journal & Thrust Bearing oil drain line was checked and found free from rust/dirt. The bearing assembly was inspected by BHEL,Hyderabad technicians. The thrust collar and end face were machined by 0.25 mm each and thrust bearing was replaced with new one. The booster pump float (6.3 mm) and thrust float (0.4 mm) were checked. BFP was then started and it again tripped on BP Thrust Bearing Temperature V.High at 108 C. Further suspecting problem in pump internal, Booster Pump Top casing was opened and impeller lock nut was checked and found in position and tight. Water Jackets were cleaned and the Booster Pump casing was assembled back. As no problems noticed in the pump, on 28.07.10 Motor was decoupled and run for about 5 minutes and the following points were observed and reported to BHEL Hyderabad and BHEL Bhopal. The rotor is moving to and fro axially to the tune of 5 mm for the initial 30 seconds. Afterwards, the shaft runs continuously in a settled position (magnetic centre) which is about 1.5 mm towards Booster pump side, from the mechanical centre. The total mechanical float of the motor rotor is found to be 12mm. Referring to the mechanical float, the motor was kept in position with respect to hydraulic coupling at one end and booster pump on the other end. Due to the offset of magnetic centre, on running the entire BFP set it is observed that the NDE side thrust pads of Booster pump is axially loaded and is getting tripped on High temperature. On 30.07.10, as suggested by BHEL Hyderabad and Bhopal Motor was shifted towards Hydraulic Coupling by 1.5 mm to match its magnetic centre. Motor alignment was done. BFP 8A was started after correcting the motor position and realignment. It tripped again on BP Thrust Bearing Temperature V. High. But this time it tripped on the DE side temperature of Thrust Bearing V. High (122.3 & 113.2 C). It was decided as per instruction of BHEL,Hyderabad to run the Pump after incorporating time delay of 5 minutes for the thrust bearing protection for observing the performance. MDBFP-A was run on 2nd Aug 2010 for about 4 minutes and following observations were made :
TECHNICAL SERVICES / PSSR

13

AUGUST 2010

TS TIDINGS

1. BP thrust pad temperature at DE side had gone Maximum value of 120.9 deg C and was coming down to 95 deg C. Thrust pad temperature at NDE was steady at around 55 deg C. 2. Movement of BP shaft was noticed back & forth during running of pump. Movement of BP shaft during running of Pump is reported to be normal & does occur during low flows; this has been reported from other sites as well. It was adviced to check flow through pump and loading of the pump. On 04.08.10 BFP was run for about 4 minutes after commissioning of flow transmitters. The following parameters were observed. Scoop position % 25% 40% Flow in TPH 122.1 & 122.5 153.6 & 154.7 BP Thrust Brg Temp DE side in deg C 113.1 & 114.9 100.8 & 89.2 BP Thrust Brg Temp NDE side in deg C 48.7 & 51.8 120.7 & 129.3

All Temperatures mentioned above are maximum temperatures during running. It was observed that BP Thrust Brg Temp DE side was fluctuating between 75 to 100.8 C. NDE side Temp was fluctuating between 70 to 129.3 C. Resolution As the temperature raise was observed even after increasing the suction flow, the thrust bearing assembly was thoroughly inspected by BHEL, Hyderabad engineers again. The following deviations were found in BFP 8A 1. Uneven thickness of BP Thrust Collar. 2. Bulging of shaft material at the collar shoulder (0.3 mm feeler gauge was going in easily between Thrust Collar and Shaft Shoulder) 3. Anti rotation pin on the floating seal had got broken. The Thrust collar was replaced with new one. Bulging was removed at the collar shoulder. Floating seal was replaced with new one. DBCF and thrust bearing clearance was checked and were found as per drawing 105.6 mm and 0.4 mm. Conclusion: After rectification of the above, the pump was started and found normal. 8 Hrs trial run of BFP 8A was done successfully and all parameters and vibrations were found to be normal. The maximum BP Thrust Bearing Temperature was around 58 Deg C.
TECHNICAL SERVICES / PSSR

14

AUGUST 2010

TS TIDINGS

Figure 1. Bearing temperature trend

TECHNICAL SERVICES / PSSR

15

AUGUST 2010

TS TIDINGS

Thrust collar With scorings

Damaged Thrust pads

Motor rotor Magnetic centre marking

TECHNICAL SERVICES / PSSR